OpenStudy (anonymous):

three times the first of three consecutive odd integers is 3 more than twice the third.What is the third integer?

hartnn (hartnn):

could u form an equation ?

OpenStudy (anonymous):

yup

hartnn (hartnn):

where u stuck ?

OpenStudy (mayankdevnani):

SORRY!! for bothering... @hartnn if x is first number, next two are x+2, and x+4 (increase of 2 each time!) so equation becomes 3x = 2(x+4)+3 solving x = 11 so numbers are x 11,13,15 three times first = 33 33 is 3 more than twice the third (2*15) = 30) yes 33 is 3 more than 30! http://answers.yahoo.com/question/index?qid=20110103190852AAqAFL2

OpenStudy (anonymous):

3x+x+1+x+3=3+2(x+3)

hartnn (hartnn):

plz don't copy answers from yahoo.com....

OpenStudy (mayankdevnani):

i don't copy them i also mention link...sir

OpenStudy (mayankdevnani):

ok @laraib

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@mayankdevnani it says odd integer

hartnn (hartnn):

wait, how did u get left side ?

OpenStudy (mayankdevnani):

yaa it gets 33 which is an odd no. @laraib

hartnn (hartnn):

if smallest is x then others are x+2.,x+4

hartnn (hartnn):

u took, x+1 and x+3 thats incorrect

hartnn (hartnn):

three times the first of three consecutive odd integers -->3x 3 more than twice the third---> 2(x+4) + 3

OpenStudy (anonymous):

What if we take (x-2), x and (x+2) ??

hartnn (hartnn):

will do.

hartnn (hartnn):

we get same answer.

OpenStudy (anonymous):

Isn't it easier one??

OpenStudy (anonymous):

3(x-2) = 2(x+2) + 3..

hartnn (hartnn):

can't say easier....just another method.

hartnn (hartnn):

both equations are equally easy to solve..

OpenStudy (anonymous):

x = 10 + 3